truck moves along the X- axis. It is at x1 = 277 m at ti= 16 sec and is at X2 = 19 m at t2= 25 Sec. Find the truck's average velocity.
1
Expert's answer
2020-11-16T07:48:58-0500

By definition, the average velocity is the total distance divided by the total time of travelling. Thus, obtain:


"v = \dfrac{\Delta x}{\Delta t}"

where "\Delta x = x_2-x_1 = 19m-277m = -258m" and "\Delta t = 25s-16s = 9s".

Substituting to the formula, get


"v = \dfrac{\Delta x}{\Delta t} = \dfrac{-258m}{9s} \approx -28.67m/s"

Answer. -28.67 m/s.